# Instruction-Guided Scene Text Recognition

### v2

@@ Abstract
Multi-modal models show appealing performance in visual recognition tasks recently, as free-form text-guided training evokes the ability to understand fine-grained visual content. However, current models are either inefficient or cannot be trivially upgraded to scene text recognition (STR) due to the composition difference between natural and text images. We propose a novel instruction-guided scene text recognition (IGTR) paradigm that formulates STR as an instruction learning problem and understands text images by predicting character attributes, e.g., character frequency, position, etc. IGTR first devises $\left\langle condition,question,answer\right\rangle$ instruction triplets, providing rich and diverse descriptions of character attributes. To effectively learn these attributes through question-answering, IGTR develops lightweight instruction encoder, cross-modal feature fusion module and multi-task answer head, which guides nuanced text image understanding. Furthermore, IGTR realizes different recognition pipelines simply by using different instructions, enabling a character-understanding-based text reasoning paradigm that considerably differs from current methods. Experiments on English and Chinese benchmarks show that IGTR outperforms existing models by significant margins, while maintaining a small model size and efficient inference speed. Moreover, by adjusting the sampling of instructions, IGTR offers an elegant way to tackle the recognition of both rarely appearing and morphologically similar characters, which were previous challenges. Code: https://github.com/Topdu/OpenOCR.
@@ 1 Introduction
Similarly, STR models could benefit significantly from an enhanced ability to understand text images thoroughly. However, given the distinct compositional elements of natural and text images, achieving this presents considerable challenges when trivially applying existing instruction learning schemes [20, 13, 14, 21, 22] to STR. Unlike natural images, which can depict objects and scenes from the physical world and be effectively described using natural language, the text image is typically described by a single word. When the word is simply treated as language, it provides limited semantic context to guide the STR model understanding the text image. Therefore, developing rich instructional strategies for STR tasks is an urgent necessity. Moreover, the text image is composed of sequentially arranged characters, necessitating the exploration of character-central properties to facilitate a deep understanding of the text. However, how to explore these properties and how they benefit text recognition are less studied in existing solutions. Furthermore, with the popularity of using OCR models in mobile and edge scenarios, where lightweight STR models [1, 5, 23] are required, it is yet to explore further multi-modal STR models that run fast and are less demanding in computational resources.
In this paper, we aim to build an efficient instruction-learning-based paradigm for STR. Our proposed method, termed instruction-guided STR (IGTR), has made meaningful attempts in both instruction construction and learning architecture design. For instruction construction, to evoke a human-like understanding and recognition of the text image, we first define character attributes as a set of statuses regarding the status, position, and frequency of character or character combinations within a text. Then, we construct two types of instructions tailored to character attribute prediction and text recognition, respectively. Specifically, we argue that character attributes are critical components to establishing a deep understanding of the text. As a result, rich and diverse character attribute prediction instructions in the form of $\left\langle condition,question,answer\right\rangle$ are devised, where condition denotes attributes already known, and question-answer is question-answer pair regarding learning the rest attributes, as shown in Tab. I. Meanwhile, recognition instructions of the same form are proposed to simulate different text recognition pipelines, including PR, AR and other useful but less-explored ones. Each can be realized simply by using one of the instructions, as shown in Tab. II.
@@
1. Char {$c_{i}$} in the image 2. Char {$c_{i}$} appear {$N_{i}$} times 3. The $i$-th char are $c_{i}$ 4. Sub-string {$c_{s}$ $c_{s+m}$} in the image 5. There are L chars in the image 6. None 1/2/3/4/5/6 How many A in the image? 2 Frequency
@@
None What are the $1$st to L-th chars in the image? $c_{1}$ $c_{L}$ Parallel recognition
$1$st to $i$-th chars are $c_{1}$ $c_{i}$ What is the $i$+$1$-th char? $c_{i+1}$ Auto-regressive recognition
$1$st to $i$-$1$-th, $i$+$1$-th to $j$-th chars are known What is the $i$-th char? $c_{i}$ Re-identification
@@
There is a sub-string $c_{i}$ $c_{j}$ What is the previous/next char of the sub-string? $c_{i-1}$/$c_{j+1}$ Extrapolating recognition
@@
With the instructions, we develop a dedicated learning architecture for both attribute prediction and text recognition. It consists of an instruction encoder to generate textual embeddings based on instructions, a cross-modal feature fusion module for characterizing interactions between image and text modalities, and a multi-task answer head responsible for answering different questions and reading the text. IGTR adopts a two-stage training process. First, it is pre-trained using a large number of attribute prediction instructions, which endows the model with a profound understanding of character attributes. Then, it is fine-tuned by recognition instructions, which enables the model’s superior text recognition capability. All attribute prediction instructions are sampled online during training and the whole architecture enjoys an end-to-end optimization.
@@
We conduct extensive experiments on English and Chinese benchmarks. The results show that IGTR has quite impressive performance. No matter following PR or AR pipeline, IGTR outperforms existing models by clear margins in terms of accuracy, while still maintaining a small model size and efficient inference speed. Meanwhile, the attribute prediction instructions guide meaningful text understanding, even when solely pre-trained on them, the model exhibits impressive zero-shot recognition capability. In addition, IGTR enjoys remarkable learning flexibility inherited from its instruction-guided nature. By simply adjusting the rule of instruction sampling, IGTR exhibits superior performance in recognizing both rarely appearing and morphologically similar characters, which were common challenges for previous methods.
@@
IGTR represents a novel paradigm of STR. As exemplified in Fig. 1, the top half represents the recognition pipeline of popular attention-based STR models. They utilize the attention mechanism to position visual features for every character and then use the features to perform character classification. Mis-recognitions may happen if features are positioned incorrectly. In contrast, as shown in the bottom half, IGTR successfully grasps character attributes and the text image by exhaustive question-answering-based learning. Therefore visual features make a correct response and IGTR generates more robust recognition. On the other hand, recent large multi-modal models such as LLaVA [21] and Monkeys [24, 22], trained or fine-tuned on OCR-relevant dialog datasets [25], also show remarkable performance in multiple OCR-related tasks including STR. However, these models typically have huge parameters and are highly demanding in computational resources. In contrast, the size of IGTR is 24.1M, and different IGTR models consume 4ms-10ms only when inferring a text instance in one NVIDIA 1080Ti GPU, both are appealing properties in mobile and edge computing applications.
@@
  • We propose IGTR that regards STR as a cross-modal instruction learning task. Unlike existing models, IGTR facilitates text recognition by comprehending text images, presenting a novel paradigm for STR.
  • We introduce rich and diverse instruction triplets $\left\langle condition,question,answer\right\rangle$, and develop a dedicated architecture to effectively learn these instructions and the text image, establishing the first instruction-guided STR solution.
  • Through extensive experiments, we demonstrate the effectiveness of IGTR not only in public benchmarks, but also in providing a uniform framework to tackle common STR challenges flexibly.
@@ 2.1 Scene Text Recognition (STR)
We can broadly classify existing sequence-based STR models into two types according to their recognition pipelines, i.e., PR and AR. PR identifies all the characters within the text image at once. It can be further classified as CTC-based and encoder-decoder-based parallel decoding ones. The former [1, 26, 5] assumes that the text image can be implicitly split into a series of stripes from left to right, each corresponding to a recognition unit, i.e., a character or blank. The recognized unit sequence is then optimized by the CTC rule [27] to get recognition results. The second [2, 28, 3, 29, 6, 7] commonly adopt an attention-based encoder-decoder framework. They first allocate a fixed number of placeholders, each corresponding to a character, and then learn how to fill in the placeholders with proper features and deduce the characters in parallel. These methods often have fast inference speed. However, early studies may not perform well when encountering text deformation. Recent studies propose various ways of utilizing language and other clues, for example, knowledge distillation [28], external language models [3, 29, 7], character counting [6], etc., which improve the recognition accuracy remarkably.
In contrast, AR [8, 9, 10, 30] adopts a one-by-one sequential recognition pipeline. It is introduced to STR following the success in NLP [31]. These models utilize previously decoded characters as language clues to assist recognition, thus getting better results in general. However, their inference speed is slow, and also may suffer from the problem of attention drift when encountering rarely seen text images like highly deformed text [32]. To mitigate attention drift, some methods [33, 32, 34] proposed to utilize the position clue. Meanwhile, permuted and bidirectional recognition are developed [11, 35], providing a stronger language prior. Generally, AR provides more natural and finer language integration compared to PR. However, only a few previously identified characters are incorporated as guidance, which limits the assistance from the language.
In addition to the two types above, there also are a few studies employing double-check techniques to assist recognition. For example, character re-identification has been considered in several methods [2, 3, 11, 36]. Typically, they use language models to correct initial results from PR or AR. Despite improving the accuracy, however, miscorrections may also happen especially for contextless text.
@@ 2.2 Multi-Modal Visual Recognition
Multi-modal models [16, 17] has received widespread attention recently, especially in visual recognition tasks, as instruction learning can evoke an appealing ability to understand fine-grained visual content. CLIP [20] and ALIGN [37] are pioneering studies in this direction. They have performed cross-modal contrastive learning on billion-level image-text pairs, enabling impressive zero-short image classification [19]. Later, GLIP [12] and Grounding-DINO [13] introduce instruction-based object detection [18]. They utilize instructions of free-form language for training, and can detect objects that exactly match the language. Similar approaches are explored in image segmentation, for instance, the well-regarded SAM [14] and SegGPT [15]. These models leverage rich interactive instructions such as clicks, boxes, masks and text to achieve generic and deep visual understanding, enabling segmentation of anything. Instruction-based recognition has also been applied to tasks like license plate recognition [38] and speech recognition [39]. However, they are more like applications of VQA models [40, 41] or generative models [42, 43] dedicated to these tasks.
Our study is also related to recent research efforts on large multi-modal models. One is prompt engineering on pre-trained visual-language models [44], which performs recognition based solely on prompts without updating model parameters. Typically, it uses an image and a text question to query the model and get the answer [45, 46], which can also reply STR-related questions. Another is large multi-modal models fine-tuned on OCR, for example, LLaVA [21] and Monkeys [24, 22]. They are capable of performing multiple OCR-related tasks like document understanding or key information extraction without relying on traditional OCR tools [23]. However, these models typically consume substantial resources in downstream applications, restricting their use in mobile and edge computing scenarios where lightweight STR models [1, 5, 23, 47] are adopted widely. To date, instruction-based STR is not well explored especially its efficient implementation.
@@ 3 Method
Our method consists of two parts, i.e., instruction construction and sampling, IGTR architecture and learning. The instructions, including both character attributes prediction and recognition ones, are $\left\langle condition,question,answer\right\rangle$ triplets. The triplet composition ensures that many instructions can be generated for attribute learning even for a short text, and meanwhile, different recognition pipelines can be precisely described. On the other hand, IGTR architecture is developed to enable efficient encoding of the instructions, and then effective cross-modal learning and recognition. They are detailed as follows.
@@ 3.1 Instruction Construction and Sampling
We mainly introduce character attribute prediction instructions here and recognition instructions will be outlined later. As shown in Tab. I, a series of character attribute prediction instructions are proposed. Each instruction consists of a condition, a question, and an answer. The condition means character attributes given in advance, and the question-answer pair corresponds to the specified query and its response.
We have meticulously crafted various instructions to ensure that the text and characters can be thoroughly described. As seen in Tab. I, the questions can be grouped into four types according to their answers (highlighted by different colors), i.e., char, frequency, position, and status. These questions compel the model to assess the relevance of each character from different aspects, determining whether it contributes to the answer. This process, in turn, guides the model towards attribute understanding.
Given a text, we further define five types of character attributes: character-status (cs) represents whether a specific character exists, character-frequency (cf) denotes the occurrence times of a specific character, constrained character-frequency (cfcons) describes character-frequency given a specific constraint (e.g., the first 3 characters), position-character (pc) means the character at the i-th position, and sub-string (ss) denotes a certain length sub-string. For example, given text ARTETA, assuming the constraint is the first 3 characters and the sub-string length $m$ is set to 3, character attributes can be obtained by traversing the text, as expressed in the following set form:
@@
Note that each element in those sets is represented by a square bracket with two units, corresponding to the variables in question and answer, and shown in Tab. I.
We propose to enlarge the traditional question-answer instruction to $\left\langle condition,question,answer\right\rangle$ triplet. Specifically, a partitioning strategy is proposed to split character attributes into two parts: one is condition to represent already known attributes, and the other is attributes left for question-answer-based model learning. Assuming that the condition is [[E,1], [R,1], [S,0]], [[E,1],[T,2]] and [[0,A], [2,T], [5,A]], which are subsets of ce, cf and pc, respectively. The condition can be described as follows: “The characters E and R exist in the image, but S does not exist”; “The character T appears twice and E appears once in the image”; and “The 1st/3rd/6-th character in the image is A/T/A”. Take the set split on pc as an example, the remaining subset is [[1,R], [3,E], [4,T]], allowing questions such as “Which character in the image is $R$/$E$/$T$?” with answers “the 2nd/4-th/5-th one” for model training. Other question-answer pairs are generated similarly. According to the types of answer, questions and answers are categorized into four types, i.e., Ques = [$Q_{c}$,$Q_{f}$,$Q_{p}$,$Q_{s}$] and Ans = [$A_{c}$,$A_{f}$,$A_{p}$,$A_{s}$], where the suffixes $c$, $f$, $p$, and $s$ denote the answer type is char, frequency, position, and status, respectively.
The triplet form notably increases the richness and diversity of instructions. To sample instructions for each input text during model training, we repeat the partitioning strategy (GenInsts) K times, with the condition randomly shuffling and splitting each time to ensure disjoint and varied instructions. This process is carried out by RandomSplit(Attributes), as described in Algorithm 1. For each partition, we enumerate all possible question-answer variables and up to $\sum\limits_{p=1}^{N}{{(L-p)}\frac{N!}{p!(N-p)!}}$ different instructions could be sampled, where $N$ is the number of attributes (len(atrb)), p represents the number of attributes in condition and it is randomly obtained using random.randint(len(atrb)). Moreover, by sampling K times, we get a large number of instructions. Consequently, the instructions and text image constitute an overwhelming many-to-one correspondence, where the language modality has been significantly enriched.

### v3

@@ Abstract
Multi-modal models have shown appealing performance in visual recognition tasks, as free-form text-guided training evokes the ability to understand fine-grained visual content. However, current models cannot be trivially applied to scene text recognition (STR) due to the compositional difference between natural and text images. We propose a novel instruction-guided scene text recognition (IGTR) paradigm that formulates STR as an instruction learning problem and understands text images by predicting character attributes, e.g., character frequency, position, etc. IGTR first devises $\left\langle condition,question,answer\right\rangle$ instruction triplets, providing rich and diverse descriptions of character attributes. To effectively learn these attributes through question-answering, IGTR develops a lightweight instruction encoder, a cross-modal feature fusion module and a multi-task answer head, which guides nuanced text image understanding. Furthermore, IGTR realizes different recognition pipelines simply by using different instructions, enabling a character-understanding-based text reasoning paradigm that differs from current methods considerably. Experiments on English and Chinese benchmarks show that IGTR outperforms existing models by significant margins, while maintaining a small model size and fast inference speed. Moreover, by adjusting the sampling of instructions, IGTR offers an elegant way to tackle the recognition of rarely appearing and morphologically similar characters, which were previous challenges. Code: https://github.com/Topdu/OpenOCR.
@@ 1 Introduction
Similarly, STR models could benefit significantly from an enhanced understanding of text images. However, given the distinct compositional elements of natural and text images, applying existing instruction learning schemes [20, 13, 14, 21, 22] to this context poses substantial challenges. Unlike natural images which primarily depict objects and scenes from the physical world and can be effectively described using natural language, text images typically represent single words. Treating a word simply as the language provides limited semantic context to guide STR models in understanding the text images. Therefore, developing dedicated instructional strategies for STR tasks is an urgent necessity. Note that the text image is composed of sequentially arranged characters, necessitating the exploration of character-central properties to facilitate a deep understanding of the text content. However, the exploration of these properties and their contributions to text recognition has received limited attention in existing solutions. Furthermore, with the increasing use of Optical Character Recognition (OCR) models in mobile and edge computing scenarios, where lightweight STR models [1, 5, 23] are required, it is yet to explore further multi-modal STR models that run fast and are less demanding in computational resources.
In this paper, we aim to build an efficient instruction-learning-based paradigm for STR. Our proposed method, termed instruction-guided STR (IGTR), firstly attempts to make meaningful efforts in both instruction construction and architecture design. In terms of instruction construction, we define character attributes as properties regarding the status, frequency, and position of one or multiple characters within text. We argue that character attributes are critical components to establishing a deep understanding of text. Then, we develop two types of instructions customized for character attribute prediction and text recognition. For character attribute prediction, we devise rich and diverse instructions in the form of $\left\langle condition,question,answer\right\rangle$, as shown in Tab. I. The condition denotes known attributes, and question-answer is question-answer pairs regarding attributes or attribute combinations. They represent learning data and labels for model training, aiming to evoke a human-like understanding of text images. Meanwhile, recognition instructions of the same form are proposed to consolidate the model training and perform text recognition. They simulate different text recognition pipelines, including PR, AR, and other useful but less-explored ones, as shown in Tab. II.
@@
1. Char {$c_{i}$} in the image 2. Char {$c_{i}$} appear {$N_{i}$} times 3. The $i$-th char are $c_{i}$ 4. Sub-string {$c_{i}$ $c_{i+l_{s}}$} in the image 5. There are L chars in the image 6. None 1/2/3/4/5/6 How many A in the image? 2 Frequency
@@
None What are the $1$st to L-th chars in the image? $c_{1}$ $c_{L}$ Parallel Recognition (PR)
$1$st to $i$-th chars are $c_{1}$ $c_{i}$ What is the $i$+$1$-th char? $c_{i+1}$ Auto-regressive Recognition (AR)
$1$st to $i$-$1$-th, $i$+$1$-th to $j$-th chars are known What is the $i$-th char? $c_{i}$ Re-Identification (RI)
@@
There is a sub-string $c_{i}$ $c_{i+l_{s}}$ What is the previous/next char of the sub-string? $c_{i-1}$/$c_{i+l_{s}+1}$ Extrapolating Recognition (ER)
@@
With the instructions, we develop a dedicated architecture for both attribute prediction and text recognition. It consists of a lightweight instruction encoder to generate textual embeddings based on instructions, a cross-modal feature fusion module for characterizing interactions between image and text modalities, and a multi-task answer head responsible for answering different questions and reading the text. IGTR is trained using a large number of attribute prediction and recognition instructions, which endows the model with a profound understanding of character attributes. Then, text images can be accurately recognized with the distinct recognition pipeline in Tab. II.
@@
We conduct extensive experiments on English and Chinese benchmarks to analyze the performance of IGTR. No matter following PR or AR pipeline, IGTR outperforms existing models by clear margins in terms of accuracy, while maintaining a small model size and fast inference speed. Meanwhile, the attribute prediction instructions guide meaningful text understanding, even when solely trained on them, IGTR exhibits impressive recognition capability. Furthermore, IGTR benefits from remarkable learning flexibility inherited from its instruction-guided nature. By simply adjusting the rule of instruction sampling, IGTR exhibits superior performance in recognizing both rarely appearing and morphologically similar characters, which have been persistent challenges for previous methods.
@@
IGTR holds the potential to represent a novel paradigm of STR. As exemplified in Fig. 1, the top half represents the recognition pipeline of popular attention-based STR models. They utilize the attention mechanism to position visual features for every character and then use the features to perform character classification. Mis-recognitions may happen if features are positioned incorrectly. In contrast, as shown in the bottom half, IGTR successfully grasps character attributes by exhaustive question-answering-based learning, and then comprehends the text image. Therefore visual features associate characters correctly and IGTR generates more robust recognition. On the other hand, recent Multi-modal Large Language Models (MLLMs) such as LLaVA [21] and Monkeys [24, 22], trained or fine-tuned on OCR-relevant dialog datasets [25], also show remarkable performance in multiple OCR-related tasks including STR. However, these models typically use large language models to compile the text modality, thus having huge parameters and being highly demanding in computational resources. In contrast, the model size of IGTR is 24.1M, and different IGTR models consume 4 ms-10 ms only when inferring a text instance in one NVIDIA 1080Ti GPU, both are appealing properties in mobile and edge computing applications.
@@
  • We propose IGTR that regards STR as a cross-modal instruction learning task. Unlike existing models, IGTR facilitates text recognition by comprehending text attributes, presenting a novel paradigm for STR.
  • We introduce rich and diverse triplet-form instructions $\left\langle condition,question,answer\right\rangle$, and develop a dedicated architecture to effectively learn these instructions and the text image, establishing the first instruction-guided STR solution.
  • Through extensive English and Chinese experiments, we demonstrate the effectiveness of IGTR not only in public benchmarks, but also in offering a versatile framework to tackle prevalent STR challenges.
@@ 2.1 Scene Text Recognition (STR)
We can broadly classify existing sequence-based STR models into two types according to their recognition pipelines, i.e., PR and AR. PR identifies all the characters within a text image at once. It can be further classified as Connectionist Temporal Classification (CTC)-based and parallel decoding ones. The former [1, 26, 5] assumes that the text image can be implicitly split into a series of stripes from left to right, each corresponding to a recognition unit, i.e., a character or blank. The recognized unit sequence is then optimized by the CTC rule [27] to get recognition results. The second [2, 28, 3, 29, 6, 7] commonly adopt an attention-based encoder-decoder framework. They first allocate a fixed number of placeholders and then learn to fill in the placeholders with proper features and deduce the characters in parallel. To compensate for semantic context, recent studies propose various ways, for example, knowledge distillation [28], external language models [3, 29, 7], character counting [6], etc., which improve the recognition accuracy remarkably.
In contrast, AR [8, 9, 10, 30] adopts a one-by-one recognition pipeline. These models leverage not only visual features but also previously decoded characters as language clues to enhance recognition accuracy, typically yielding superior results. However, they may encounter challenges such as attention drift when faced with rarely seen patterns like highly deformed text [31]. To mitigate attention drift, some methods propose to incorporate the position clue into the recognition process [32, 31, 33]. Additionally, methods like permuted and bidirectional recognition are developed [11, 34], providing a stronger language prior. Generally, AR models provide more natural and finer language integration compared to PR models. However, their sequential decoding nature results in slower inference speed, limiting their use in time-sensitive applications.
In addition to the two types above, there are studies employing double-check-like techniques to assist recognition. For example, some methods incorporate character re-identification strategies [2, 3, 11, 35]. Typically, they use language models to rectify initial outputs generated by PR or AR models. Despite improving the accuracy, they also incur a possibility of introducing errors, especially when dealing with contextless text.
@@ 2.2 Multi-Modal Visual Recognition
Multi-modal models [16, 17] have achieved promising performance recently, particularly in visual recognition tasks, as instruction learning can evoke an appealing understanding of fine-grained visual content. CLIP [20] and ALIGN [36] are pioneering studies in this direction. They have performed cross-modal contrastive learning on billion-level image-text pairs, enabling impressive zero-short image classification [19]. Later, GLIP [12] and Grounding-DINO [13] introduce instruction-based object detection [18]. They utilize instructions of free-form language for training, and can detect objects that exactly match the language. Similar approaches are explored in image segmentation, for instance, the well-regarded SAM [14] and SegGPT [15]. These models leverage rich interactive instructions including text to achieve generic and deep visual understanding, enabling segmentation of anything. Generally, these studies mainly focused on more generic visual tasks which involve multi-object, multi-scene contexts. However, text images primarily focus on character-central information. This distinction requires a specific adaptation of the instruction-guided approach to capture the fine-grained attributes unique to text, which is seldom discussed before. Note that instruction-based recognition has also been applied to tasks like license plate recognition [37] and speech recognition [38]. However, they are more like the application of VQA models [39, 40] or generative models [41, 42] to these tasks.
Note that recent multi-modal large language models (MLLMs) [43, 21, 24, 22] have also shown impressive performance on OCR tasks. These models typically implement OCR in two ways. One branch is prompt engineering on pre-trained visual-language models [44], which performs recognition based solely on prompts without updating model parameters. Typically, this approach involves querying the model with an image and a text question to generate responses [45, 43], which can extend to handling OCR-related queries. Another branch focuses on fine-tuning MLLMs on OCR data, for example, LLaVA [21] and Monkeys [24, 22]. These models demonstrate capabilities in various OCR-related tasks like document understanding and key information extraction without relying on traditional OCR tools [23]. Nevertheless, the two branches above typically incorporate large language models to better capture textual information. As a result, they consume substantial resources, restricting their deployment in mobile and edge computing scenarios where lightweight STR models [1, 5, 23, 46] are prevalent. In contrast, IGTR simplifies textual query modeling into a straightforward embedding format, eliminating the need for large language models and significantly reducing the model’s computational overhead.
@@ 3 Method
Our method consists of two parts, i.e., instruction construction and sampling, IGTR architecture and learning. The instructions, including both character attributes prediction and recognition ones, are $\left\langle condition,question,answer\right\rangle$ triplets. The triplet composition ensures that many instructions can be generated for attribute learning even for a short text, and meanwhile, different recognition pipelines can be precisely described. On the other hand, IGTR architecture mainly consists of a lightweight instruction encoder, a cross-modal feature fusion module and a multi-task answer head. They together enable efficient encoding of the instructions, and effective cross-modal learning and recognition.
@@ 3.1 Instruction Construction and Sampling
We have defined a series of character attribute prediction and text recognition instructions. Each instruction consists of a condition, a question, and an answer. The condition means knowledge given in advance, and the question-answer pair corresponds to the specified query and its response. In Tab. I, Condition Option describes the compatibility between conditions and question-answers for character attribute prediction instructions. These questions, together with the corresponding conditions, compel the model to assess the character from different aspects, determining whether it contributes to the answer. This process, in turn, guides the model towards nuanced text understanding. Note that the question-answers in Tab. I can be classified into four types according to their answers (highlighted by different colors in Type column), i.e., character, frequency, position, and status.
We further define five types of character attributes: character-status (cs) represents whether a specific character exists, character-frequency (cf) denotes the occurrence times of a specific character, constrained character-frequency (cfcons) describes character-frequency given a specific constraint (e.g., the first 3 characters), position-character (pc) means the character at the i-th position, and sub-string (ss) denotes a certain length sub-string. For example, given text ARTETA, assuming the constraint is the first 3 characters and the sub-string length $l_{s}$ is set to 3, character attributes can be obtained by traversing the text, as expressed in the following set form:
@@
Note that each element in those sets is represented by a square bracket with two units, corresponding to the variables in question and answer, respectively.
We propose to enlarge the traditional question-answer instruction to $\left\langle condition,question,answer\right\rangle$ triplet. Specifically, a partitioning strategy is proposed to split character attributes into two parts: one is condition to represent already known attributes, and the other is attributes left for question-answer-based model learning. Assuming that the condition is [[E,1], [R,1], [S,0]], [[E,1],[T,2]] and [[0,A], [2,T], [5,A]], which are subsets of ce, cf and pc, respectively. The condition can be described as follows: “Characters E and R exist in the image, but S does not exist”; “Character T appears twice and E appears once in the image”; and “The 1st/3rd/6-th character in the image is A/T/A”. Take the set split on pc as an example, the remaining subset is [[1,R], [3,E], [4,T]], allowing questions such as “Which character in the image is $R$/$E$/$T$?” with answers “the 2nd/4-th/5-th one” for model training. Other question-answer pairs are generated similarly. According to the types of answer, questions and answers are categorized into four types, i.e., Ques = [$Q_{c}$,$Q_{f}$,$Q_{p}$,$Q_{s}$] and Ans = [$A_{c}$,$A_{f}$,$A_{p}$,$A_{s}$], where the suffixes $c$, $f$, $p$, and $s$ denote the answer type is character, frequency, position, and status, respectively.
The triplet form notably increases the richness and diversity of instructions. To sample instructions for each input text during model training, we repeat the partitioning strategy (GenInsts) K times, with the condition randomly shuffling and splitting each time to ensure disjoint and varied instructions. This process is carried out by RandomSplit(Attributes), as described in Algorithm 1. For each partition, we enumerate all possible question-answer variables and up to $\sum\limits_{p=1}^{N}{{(L-p)}\frac{N!}{p!(N-p)!}}$ different instructions could be sampled, where $N$ is the number of attributes len(atrb), p represents the number of attributes in condition and it is randomly obtained using random.randint(len(atrb)). By sampling K times, we can get an even larger number of instructions.